# Direct and inverse abstract Cauchy problems with fractional powers of almost sectorial operators

### v1

@@ 4. Inverse abstract Cauchy problems
We now use the representation of the solution operator of the (direct) abstract Cauchy problem (found in the previous Section) to find the time-dependent coefficient for the inverse abstract Cauchy problem. The method is based in the consideration of two (direct) abstract Cauchy problems with the same time-dependent coefficient.
We start by establishing a useful property that involves the analytic semigroup of growth order and a fractional power of an operator. Let $A\in\Theta_{\omega}^{\gamma}$ for some $20<\omega<\pi/$ 2 . From [14, Theorem 2.5] and Theorem 2.3 (item (3)), we get that
@@
since the function $z^{\alpha}e^{-tz^{\alpha}}$ is in the class $\mathcal{F}_{0}^{\gamma}(S_{\mu}^{0})$ $\big{(}\omega<\theta<\mu<\text{min}\left\{\pi,\frac{\pi}{2\alpha}\right\}\big{)}.$ More details about it can be found in [14, Section 2].
@@
(4.2) $\begin{split}u_{t}(t)&=\phi(t)A^{\alpha}u(t),\quad 0<t<T,\quad 0<\alpha<\frac{\pi}{2\omega},\\ u(0)&=u_{0}\in D(A^{\alpha}),\end{split}$
@@
(4.3) $\begin{split}v_{t}(t)&=\phi(t)A^{\alpha}v(t),\quad 0<t<T,\quad 0<\alpha<\frac{\pi}{2\omega},\\ v(0)&=A^{\alpha}u_{0}\in D(A^{\alpha}).\end{split}$
By Theorem 3.1, the solution operators of problems (4.2) and (4.3) are given respectively by
(4.4) $u(t)=\mathscr{T}_{\alpha}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0},\quad\text{and}\quad v(t)=\mathscr{T}_{\alpha}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}A^{\alpha}u_{0}.$
@@
$h_{1}(t):=u(t,q),\quad h_{2}(t):=v(t,q),\quad 0<t<T.$
Next, we establish the main result of this section as follows.
@@
By using (4.4), Theorem 2.3 (item (4)) and property (4.1) we obtain
@@
Again, it is worthy to highlight the particular case $\alpha=1.$ Hence, let us study the following two abstract Cauchy problems with the variable coefficient $\phi$:
@@
$\widetilde{h_{1}}(t)=\mathscr{T}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0}(t,q)\quad\text{and}\quad\widetilde{h_{1}}(t)=\mathscr{T}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}Au_{0}(t,q).$
@@ 5. Non-autonomous Cauchy problems
Suppose now that our linear operator $A$ depends also of a time-variable $\tau\geqslant 0$. In this case, the operator is usually denote by $A(\tau).$ By doing an analog of the classical definition, in [3, 8], the family of all linear closed operators $A(\tau):D(A(\tau))\subset X\to X$ which satisfy both conditions of Definition 2.1 with $A(\tau)$ instead of $A$, it is called an almost sectorial operator and it is again denoted by $\Theta_{\omega}^{\gamma}(X).$
Having in mind the study made in [1, 14], in [3, 8], it is then constructed and sometimes proved similarly many basic results for the case of $A(\tau).$ In this regard, we have that the semigroup $\{T_{A(\tau)}(t)\}_{t\in S_{\frac{\pi}{2}-\omega}}^{0}$ $(\omega<\theta<\mu<\frac{\pi}{2}-|\text{arg}\,t|,\,\text{for each}\,\tau\geqslant 0)$ of growth order $1+\gamma$ is given by
$T_{A(\tau)}(t)=\frac{1}{2\pi i}\int_{\Gamma_{\theta}}e^{-tz}(z-A(\tau))^{-1}x\,{\rm d}z,\quad t\in S_{\frac{\pi}{2}-\omega}^{0},$
and it is a bounded linear operator in $X.$ Here, the semigroup of growth order $1+\gamma$ means: i) $T_{A(\tau)}(0)=I;$ ii) $T_{A(\tau)}(t+s)=T_{A(\tau)}(t)T_{A(\tau)}(s)$ for each $t,s>0;$ iii) $T_{A(\tau)}(t)x=0$ implies $x=0$ for $t>0;$ iv) $\lim_{t\to 0}t^{1+\gamma}T_{A(\tau)}(t)$ is bounded.
@@
(5.1) $\begin{split}u_{t}(t)&=A(\tau)u(t),\quad 0<t<T,\quad\tau\geqslant 0,\\ u(0)&=u_{0}\in X.\end{split}$
By a classical solution of problem (5.1), we mean a function $u\in C\big{(}(0,T);X\big{)}\cap C^{1}\big{(}(0,T);X\big{)}$ that takes values in $D(A(\tau))$ for any $0<t<T$ and satisfies (5.1).
It is known that (see e.g. [3, Lemma 2.3]) that equation (5.1) has a unique classical solution given by
(5.2) $u(t)=T_{A(\tau)}(t)u_{0},\quad\text{for any}\quad 0<t<T,\quad\tau\geqslant 0.$
@@ 5.1. Direct abstract Cauchy problem
We study the abstract Cauchy problem with the coefficient $\phi(t)$ and $A(\tau)\in\Theta_{\omega}^{\gamma}$ $\big{(}0<\omega<\pi/2,\,\tau\geqslant 0\big{)}$ as follows:
(5.3) $\begin{split}u_{t}(t)&=\phi(t)A(\tau)u(t),\quad 0<t<T,\quad\tau\geqslant 0,\\ u(0)&=u_{0}\in X.\end{split}$
The proof of the next statement follows the same calculations and steps of the proof of Theorem 3.1. Therefore, we omit its proof and leave it to an interested reader.
@@
Let $X$ be a complex Banach space. Then the solution operator of problem (5.3) is expressed as:
(5.4) $T_{A(\tau),\phi}(t)=T_{A(\tau)}\left\{\left(\int_{0}^{t}\phi(s){\rm d}s\right)\right\},\quad 0<t<T,\quad\tau\geqslant 0.$
@@ 5.2. Inverse abstract Cauchy problems
We start by establishing a property that involves the semigroup $T_{A(\tau),\phi}(t)$ of growth order $1+\gamma$ and the operator $A(\tau)$. Let $A(\tau)\in\Theta_{\omega}^{\gamma}$ for some $20<\omega<\pi/$ 2 and $\tau\geqslant 0.$ From [3, Pages 22-24] (or [8, Lemma 2.7 (ii)]), we know that
(5.5) $A(\tau)T_{A(\tau),\phi}(t)x=T_{A(\tau),\phi}(t)A(\tau)x,\quad t\in S^{0}_{\frac{\pi}{2}-\omega},\quad x\in D(A(\tau)),\quad\tau\geqslant 0.$
@@
(5.6) $\begin{split}u_{t}(t)&=\phi(t)A(\tau)u(t),\quad 0<t<T,\quad\tau\geqslant 0,\\ u(0)&=u_{0}\in D(A(\tau)),\end{split}$
@@
(5.7) $\begin{split}v_{t}(t)&=\phi(t)A(\tau)v(t),\quad 0<t<T,\quad\tau\geqslant 0,\\ v(0)&=A(\tau)u_{0}\in D(A(\tau)).\end{split}$
By Theorem 5.1, the solution operators of problems (5.6) and (5.7) are represented by
(5.8) $u(t)=T_{A(\tau)}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0},\,\,\text{and}\,\,v(t)=T_{A(\tau)}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}A(\tau)u_{0}.$
@@
$k_{1}(t):=u(t,q),\quad k_{2}(t):=v(t,q),\quad 0<t<T.$
Next, we establish the main result of this subsection.
@@
$\phi(t)=\frac{\partial_{t}k_{1}(t)}{k_{2}(t)},\quad\text{for any}\quad 0<t<T.$
@@
By using (5.8), [3, Lemma 2.3] (or [8, Lemma 2.7 (i)]) and property (5.5) it follows that
@@
that finishes the proof. ∎
@@
Note that equations in this section depend on operators $A(\tau)$ for some $\tau\geqslant 0.$ Therefore, we study, in principle, an evolution equation with respect to the time-variable $t$ and an almost sectorial operator $A(\tau).$ The case $A(t)$ (we mean e.g. $u_{t}(t)=A(t)u(t);$ $u(\tau)=u_{0};$ $\tau>t$) is in a sense different since the solution operator is associated to a new notion of linear evolution process introduced in [3]. Also, the case of fractional powers of $A(\tau)$ is not well studied yet. Therefore, these two cases will be analyzed somewhere else.
@@
The authors declare that there are no conflicts of interest.

### v2

@@ 4. Inverse abstract Cauchy problems
We now use the representation of the solution operator of the (direct) abstract Cauchy problem (found in the previous Section) to find the time-dependent coefficient for the inverse abstract Cauchy problem. The method is based on the consideration of two (direct) abstract Cauchy problems with the same time-dependent coefficient.
We start by establishing a useful property that involves the analytic semigroup of growth order and the fractional power of an operator. Let $A\in\Theta_{\omega}^{\gamma}$ for some $0<\omega<\pi/2.$ From [16, Theorem 2.5] and Theorem 2.4 (item (3)), we have
@@
since the function $z^{\alpha}e^{-tz^{\alpha}}$ is in the class $\mathcal{F}_{0}^{\gamma}(S_{\mu}^{0})$ $\big(\omega<\theta<\mu<\text{min}\left\{\pi,\frac{\pi}{2\alpha}\right\}\big).$ More details about it can be found in [16, Section 2].
@@
(4.2) $\begin{split}u_{t}(t)&=\phi(t)A^{\alpha}u(t),\quad 0<t<T,\quad 1+\gamma<\alpha<\frac{\pi}{2\omega},\\ u(0)&=u_{0}\in D(A^{\alpha}),\end{split}$
@@
(4.3) $\begin{split}v_{t}(t)&=\phi(t)A^{\alpha}v(t),\quad 0<t<T,\quad 1+\gamma<\alpha<\frac{\pi}{2\omega},\\ v(0)&=A^{\alpha}u_{0}\in D(A^{\alpha}).\end{split}$
From now on, we always assume $u_{0}\neq 0$ (respectively $Au_{0}\neq 0$). Note that if $u_{0}=0$ then the classical solution $u$ of (4.2) (respectively (4.3)) satisfies $u\equiv 0$ for $0<t<T.$ These latter cases are not considered to recover the variable coefficient $\phi$ above.
By Theorem 3.1, the solutions of problems (4.2) and (4.3) are given, respectively, by
(4.4) $u(t)=\mathscr{T}_{\alpha}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0},\,\text{and}\,v(t)=\mathscr{T}_{\alpha}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}A^{\alpha}u_{0}.$
@@
$h_{1}(t):=u(t,q)\neq 0,\quad h_{2}(t):=v(t,q)\neq 0,\quad 0<t<T.$
This point is not an element itself of the Banach space $X.$ In fact, the point belongs to the measure space where we are defining $X.$ For instance, consider the Banach space $C^{l}(\overline{\Omega})$ for $0<l<1$ and $\Omega$ a bounded domain in $\mathbb{R}^{n}$ $(n\geqslant 1)$ with a smooth boundary. In this case, we take $q\in\Omega.$ While, in the space $L^{3}(\mathbb{R}^{2}),$ we then take $q\in\mathbb{R}^{2}.$ In the latter spaces we can always find and define different types of almost sectorial operators, see e.g. [16, Section 2] or [22, Section 6].
Example 4.1.
Let $\Omega$ be a bounded domain in $\mathbb{R}^{n}$ $(n\geqslant 1)$ with boundary $\partial\Omega$ of class $C^{4}.$ Let $X=C^{\beta}(\Omega)$ $(0<\beta<1).$ Set
$\Delta=\sum_{k=1}^{n}\frac{\partial^{2}}{\partial x_{k}^{2}},\quad D(\Delta)=\{u\in C^{2+\beta}(\Omega):\,\,u(0)=0\,\,\text{on}\,\,\partial\Omega\}.$
From [16, Example 2.3], it follows that there exist $\nu,\rho>0,$ such that
$\Delta+\nu\in\Theta_{\frac{\pi}{2}-\rho}^{\frac{\beta}{2}-1}\big(C^{\beta}(\Omega)\big).$
Consider the following initial value problem for the heat equation on $\mathbb{R}^{n}:$
(4.5) $\begin{split}w_{t}(t,x)&=\phi(t)\big(\Delta_{x}+\nu\big)w(t,x),\quad x\in\Omega\subset\mathbb{R}^{n},\,\,t>0,\\ w(t,x)&|_{{}_{{}_{t=0+}}}=g(x),\quad x\in\Omega,\\ w(t,x)&|_{{}_{{}_{t=0+}}}=0,\quad x\in\partial\Omega,\end{split}$
where $\text{supp}(g)\subset\Omega,$ $g,\widehat{g}\in L^{1}(\mathbb{R}^{n})$. If $g\in D(\Delta)$ then there exists a classical solution of problem 4.5 (see Theorem 3.1). The solution can be also found by applying the space Fourier transform and then solving the transform equation, which is an ordinary differential equation with respect to the variable $t$. So, the solution is given by:
(4.6) $w(t,x)=\frac{e^{\nu\int_{0}^{t}\phi(s){\rm d}s}}{\left(4\pi\int_{0}^{t}\phi(s){\rm d}s\right)^{n/2}}\int_{\Omega}e^{-\frac{|x-y|^{2}}{\int_{0}^{t}\phi(s){\rm d}s}}g(y)dy,\quad t>0,\,\,x\in\Omega.$
Note that from representation (4.6) it is clear that we have many possibilities for $q\in\Omega$ such that $h_{1}(t)=w(t,q)\neq 0$ as well as $h_{2}(t)\neq 0.$
We establish the main result of this section as follows.
@@
Using (4.4), Theorem 2.4 (item (4)) and property (4.1) we obtain
@@
Again, it is worth highlighting the particular case $\alpha=1.$ Hence, let us study the following two abstract Cauchy problems with the variable coefficient $\phi$:
@@
$\widetilde{h_{1}}(t)=\mathscr{T}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0}(t,q)\,\,\text{and}\,\,\widetilde{h_{2}}(t)=\mathscr{T}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}Au_{0}(t,q).$
@@ 5. Non-autonomous Cauchy problems
In this section, we suppose that our linear operator $A$ depends also on a time-variable $\tau\geqslant 0$. In this case, the operator is usually denoted by $A(\tau).$ We will work with the following class of operators [8, Definition 2.3] (see also [3]) which we also denote by $\Theta_{\omega}^{\gamma}(X)$ because it is adapted from Definition 2.1.
Definition 5.1.
For each $\tau\geqslant 0.$ Let $1<\gamma<0$ and $0\leqslant\omega<\pi.$ By $\Theta_{\omega}^{\gamma}(X)$ we denote the set of all closed linear operators $A(\tau):D(A(\tau))\subset X\to X$ that satisfy
  1. (a)
    $\sigma(A(\tau))\subset S_{\omega}.$
  2. (b)
    For any $\omega<\mu<\pi$, there exists a positive constant $C_{\mu}$ independent of $\tau$ such that
    $\|(z-A(\tau))^{-1}\|\leqslant C_{\mu}|z|^{\gamma},\quad\text{for any}\quad z\notin S_{\mu}.$
A linear operator $A(\tau)$ will be called an almost sectorial operator in $X$ if $A(\tau)\in\Theta_{\omega}^{\gamma}(X)$ for each $\tau\geqslant 0.$ In this scenario, we can also generate a type of semigroup with $A(\tau).$ In fact, by [8, Lemma 2.7] (see also [3, Pages 634–637]), the semigroup $\{T_{A(\tau)}(t)\}_{t\in S_{\frac{\pi}{2}-\omega}}^{0}$ $(\omega<\theta<\mu<\frac{\pi}{2}-|\text{arg}\,t|,\,\text{for each}\,\tau\geqslant 0)$ of growth order $1+\gamma$ is studied and defined as follows
$T_{A(\tau)}(t)=\frac{1}{2\pi i}\int_{\Gamma_{\theta}}e^{-tz}(z-A(\tau))^{-1}\,{\rm d}z,\quad t\in S_{\frac{\pi}{2}-\omega}^{0},$
which is a bounded linear operator in $X,$ such that: i) $T_{A(\tau)}(0)=I;$ ii) $T_{A(\tau)}(t+s)=T_{A(\tau)}(t)T_{A(\tau)}(s)$ for each $t,s>0;$ iii) $T_{A(\tau)}(t)x=0$ implies $x=0$ for $t>0;$ iv) $\lim_{t\to 0}t^{1+\gamma}T_{A(\tau)}(t)$ is bounded; (v) the function $t\mapsto T_{A(\tau)}(t)$ is analytic in $S_{\frac{\pi}{2}-\omega}^{0}$ and $\frac{d^{n}}{dt^{n}}T_{A(\tau)}(t)=\big(-A(\tau)\big)^{n}T_{A(\tau)}(t)$ for all $t\in S_{\frac{\pi}{2}-\omega}^{0}.$
@@
(5.1) $\begin{split}u_{t}(t)&=A(\tau)u(t),\quad t>0,\quad\tau\geqslant 0,\\ u(0)&=u_{0}\in X.\end{split}$
By a classical solution of problem (5.1), we mean a function
$u\in C\big([0,+\infty);X\big)\cap C^{1}\big((0,+\infty);X\big)$
that takes values in $D(A(\tau))$ for any $t>0$ and satisfies (5.1). It is known (see e.g. [3, Lemma 2.3] or [8, Lemma 2.7]) that the equation (5.1) has a unique classical solution given by
(5.2) $u(t)=T_{A(\tau)}(t)u_{0},\quad\text{for any}\quad t\geqslant 0,\quad\tau\geqslant 0.$
@@ 5.1. Direct abstract Cauchy problem
We study the abstract Cauchy problem with a coefficient $\phi(t)$ and $A(\tau)\in\Theta_{\omega}^{\gamma}$ $\big(0<\omega<\pi/2,\,\tau\geqslant 0\big).$ Here, we assume that $\phi:[0,+\infty)\to[0,+\infty)$ is a continuous function such that $\phi(t)>0$ for $t>0.$ The problem to be studied reads as follows:
(5.3) $\begin{split}u_{t}(t)&=\phi(t)A(\tau)u(t),\quad t>0,\quad\tau\geqslant 0,\\ u(0)&=u_{0}\in X.\end{split}$
The proof of the next statement follows the same calculations and steps as the proof of Theorem 3.1. Therefore, we omit the proof and leave it to the interested reader.
@@
Let $X$ be a complex Banach space. Then the solution operator of the problem (5.3) is expressed as:
(5.4) $T_{A(\tau),\phi}(t)=T_{A(\tau)}\left\{\left(\int_{0}^{t}\phi(s){\rm d}s\right)\right\},\quad t\geqslant 0,\quad\tau\geqslant 0.$
@@ 5.2. Inverse abstract Cauchy problems
We start by establishing a property that involves the semigroup $T_{A(\tau),\phi}(t)$ of growth order $1+\gamma$ and the operator $A(\tau)$. Let $A(\tau)\in\Theta_{\omega}^{\gamma}$ for some $0<\omega<\pi/2$ and $\tau\geqslant 0.$ From [3, Pages 22-24] (or [8, Lemma 2.7 (ii)]), we know that
(5.5) $A(\tau)T_{A(\tau),\phi}(t)x=T_{A(\tau),\phi}(t)A(\tau)x,\quad t\in S^{0}_{\frac{\pi}{2}-\omega},\,\,x\in D(A(\tau)),\,\,\tau\geqslant 0.$
@@
(5.6) $\begin{split}u_{t}(t)&=\phi(t)A(\tau)u(t),\quad t>0,\quad\tau\geqslant 0,\\ u(0)&=u_{0}\in D(A(\tau)),\end{split}$
@@
(5.7) $\begin{split}v_{t}(t)&=\phi(t)A(\tau)v(t),\quad t>0,\quad\tau\geqslant 0,\\ v(0)&=A(\tau)u_{0}\in D(A(\tau)).\end{split}$
By Theorem 5.2, the solution operators of the problems (5.6) and (5.7) are represented by
(5.8) $u(t)=T_{A(\tau)}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}u_{0},\,v(t)=T_{A(\tau)}\left\{\left(\int_{0}^{t}\phi(r){\rm d}r\right)\right\}A(\tau)u_{0}.$
@@
$k_{1}(t):=u(t,q)\neq 0,\quad k_{2}(t):=v(t,q)\neq 0,\quad 0<t<T.$
We now provide the main result of this subsection.
@@
$\phi(t)=\frac{\partial_{t}k_{1}(t)}{k_{2}(t)},\quad\text{for any}\quad t>0.$
@@
Using (5.8), [3, Lemma 2.3] (or [8, Lemma 2.7 (i)]) and property (5.5) it follows that
@@
which finishes the proof. ∎
@@
Note that the equations in this section depend on operators $A(\tau)$ for some $\tau\geqslant 0.$ Therefore, we study, in principle, an evolution equation with respect to the time-variable $t$ and an almost sectorial operator $A(\tau),$ where $t$ and $\tau$ do not have to be related. The case $A(t)$ (we mean, e.g. $u_{t}(t)=A(t)u(t);$ $u(\tau)=u_{0};$ $\tau>t$) is totally different since the solution operator is associated with a new notion of a linear evolution process introduced in [3]. We do not study this case because this will be done elsewhere.
@@
The author declares that there are no conflicts of interest.